Treat the following parts as three independent situations.

Part 1 of 3

Lynbrook Automotive, Inc. applies the allowance method to record transactions relating to their accounts receivables. On December 31, 2019, the company had a debit balance of $700,000 in Accounts Receivable and a credit balance of $20,000 in their Allowance for Doubtful Accounts. During 2020, the company completed a number of transactions involving credit sales, accounts receivable collections, and bad debts. These transactions are summarized as follows:

  1. a) Provided $1,400,000 of delivery services on credit, terms n/30.

  2. b) Wrote off $17,000 of uncollectible accounts receivable.

  3. c) Received $2,000 from customers whose accounts had previously been written off.

  4. d) Received $1,750,000 cash in payment of accounts receivable.

  5. e) In adjusting the accounts on December 31, 2020, the company estimated that 3.5% of

    credit sales will be uncollectible.

Required:

  1. Prepare journal entries to record Lynbrook’s summarized transactions and its year-end adjustment to record bad debt expense. (Round amounts to the nearest dollar.)

  2. Determine the Net Realizable Value (NRV) reported on the Balance Sheet as of December 31, 2020.

Part 2 of 3

During the fiscal year ended December 31, 2020, Lynbrook Inc. had credit sales of $4,780,000. On December 31, 2020, Lynbrook had an unadjusted debit balance in their Allowance for Doubtful Accounts of $15,275. Your accounting department prepared an aging schedule of Lynbrook’s December 31, 2020, accounts receivable account. Based on past experience, Lynbrook estimates the percent of receivables in each age category that will become uncollectible. This information is summarized here.

December 31, 2020 Age of Expected Percent Accounts Receivable Accounts Receivable Uncollectible

`

Required

1. Estimate the required balance of the Allowance for Doubtful Accounts at December 31, 2020, using the aging of accounts receivable method.

$450,000

Not yet due

1.25%

275,000

1 to 30 days past due

5.00%

89,000

31 to 60 days past due

10.00%

42,000

61 to 90 days past due

35.75%

15,000

Over 90 days past due

71.00%

  1. Prepare the adjusting entry to record bad debt expense at December 31, 2020.

  2. On February 14, 2021, Lynbrook concludes that a customer's $4,000 receivable (created in early 2020) is uncollectible and that the account should be written off. What effect will this action have on Lynbrook’s 2021 net income? Explain.

Part 3 of 3

Lynbrook Inc. sells a variety of merchandise to retail stores on account, but it insists that any customer who fails to pay an invoice when due must replace it with an interest-bearing note. The company makes annual adjustments on December 31.

On October 1, 2020 Lynbrook converted an Accounts Receivable balance from Scan House in the amount of $20,000 to a 4 month 12% Note Receivable.

Required:

A. Prepare journal entries (in general journal form) to record:
(1) the creation of the note on October 1, 2020;
(2) the adjusting journal entry on December 31, 2020;
(3) the collection of principal and interest when it matures on February 1, 2021.

By the way no need to write detailed on the comments just because you are too lazy to do it. Thanks!
Answer #1

Part 1
Accounts Receivable as on Dec 31, 2019 = $700000+1400000-17000+2000-2000-1750000 = $333000
Allowance for Doubtful Debt unadjusted Balance = $20000 - 17000+2000 = $5000
Allowance for Doubtful Debt Balance = $333000 x 3.5% = $11655
Bad Debt Expense = $6655

Account Titles Debit Credit
a Accounts Receivable $        1,400,000
      Delivery Service Revenue $      1,400,000
(Provided delivery service on account)
b Allowance for Doubtful Debts $              17,000
      Accounts Receivable $            17,000
(Accounts receivable written off)
c. Accounts Receivable $                2,000
       Allowance for Doubtful Debts $              2,000
(Accounts receivable restated)
Cash $                2,000
     Accounts Receivable $              2,000
(Cash Collected)
d Cash $        1,750,000
     Accounts Receivable $      1,750,000
(Cash Collected)
e Bad Debt Expense $                6,655
       Allowance for Doubtful Debts $              6,655
(Bad Debt expense recorded)
Accounts Receivable $            333,000
Less Allowance for Doubtful Debts $              11,655
Net Realizable Value $          321,345

Part 2

Allowance for Doubtful Debts
Not yet due $            450,000 1.25% $                  5,625
1 to 30 days past due $            275,000 5.00% $                13,750
31 to 60 days past due $              89,000 10.00% $                  8,900
61 to 90 days past due $              42,000 35.75% $                15,015
Over 90 days past due $              15,000 71.00% $                10,650
$                53,940


Bad Debt Expense = $53940 - 15275 = $38665

Account Titles Debit Credit
Bad Debt Expense $              38,665
       Allowance for Doubtful Debts $            38,665
(Bad Debt expense recorded)

Writing off receivable will not have any direct effect on net income, but writing off receivable will reduce balance of allowance for doubtful debts as well which will increase bad debt expense at year end and hence lower net income

Part 3

Date Account Titles Debit Credit
Oct-01 Note Receivable $              20,000
       Accounts Receivable $            20,000
Dec-31 Interest Receivable $                    600 =20000*12%*3/12
      Interest Revenue $                  600
(Interest recorded for 3 months)
Feb-01 Cash $              20,800
      Note Receivable $            20,000
      Interest Receivable $                  600 =20000*12%*3/12
      Interest Revenue $                  200 =20000*12%*1/12